A population consists of four numbers 3, 7, 11, 15. consider all possible sample size 2 which can be drawn without replacement. Find mean and SD in the population.


1
Expert's answer
2021-05-23T15:21:15-0400

"Number = \frac{N!}{n!(N-n)!} \\

= \frac{4!}{2!(4-2)!} \\

= \frac{3 \times 4}{2} \\

= 6"

Samples:

3,7

3,11

3,15

7,11

7,15

11,15

"Mean = \frac{3+7+11+15}{4} = 9 \\

SD = \sqrt{\frac{1}{4-1}( (3-9)^2+(7-9)^2 +(11-9)^2 + (15-9)^2 )} \\

= \sqrt{ \frac{1}{3}(36+4+4+36) } \\

= 5.164"